# Using neuroevolution for designing soft medical devices

### v1

@@ Abstract
Soft robots can exhibit better performance in specific tasks compared to conventional robots, particularly in healthcare-related tasks. However, the field of soft robotics is still young, and designing them often involves mimicking natural organisms or relying heavily on human experts’ creativity. A formal automated design process is required. We propose the use of neuroevolution-based algorithms to automatically design initial sketches of soft actuators that can enable the movement of future medical devices, such as drug-delivering catheters. The actuator morphologies discovered by algorithms like Age-Fitness Pareto Optimization, NeuroEvolution of Augmenting Topologies (NEAT), and Hypercube-based NEAT (HyperNEAT) were compared based on the maximum displacement reached and their robustness against various control methods. Analyzing the results granted the insight that neuroevolution-based algorithms produce better-performing and more robust actuators under different control methods. Moreover, the best-performing morphologies were discovered by the NEAT algorithm. As a future work aspect, we propose using the morphologies discovered here as test beds to optimize specialized controllers, enabling more effective functionality towards the desired deflections of the suggested soft catheters.
@@ 1 Introduction
The concept of soft robots can be traced back to Isaac Asimov’s books published in the 1940s-1950s. However, the first scientific and engineering papers related to soft grippers [20] and soft architecture machines [42] started in the late 1970s. The domain has started to flourish in late 1980s with further publications on soft fingers [3, 9, 10], electrostatic actuators [16], and fluid based grippers [25]. Unlike traditional and rigid robots, these robots are built utilising flexible and ductile materials [32, 50]. Soft robots have the potential to exhibit better performance in specific tasks that are related to search and rescue [39, 41, 62, 64], space exploration [38, 5, 43, 40], and healthcare [46, 33, 22] due to their mechanisms and morphology, which are inspired by the movement and behaviour of living organisms.
Publications on soft robotic catheters, and relevant surgical devices, started to emerge from late 1980s [29] and by 2020 there was a range of catheters developed for cardio-vascular surgery [45, 49, 44], urinary system [30, 8, 65], endovascular treatment [18], in situ bioprinting [67], endoscopy [31], drug delivery [23]. Plantoid robots proposed by Mazzolai and colleagues [35, 37, 15, 36, 34, 51] can be considered relevant to the research on robotic catheters. These robots are soft-bodied mechanical devices capable of physically growing at the tip (thanks to embedded 3D printing techniques), similar to a plant root, and can change the direction of their growth due to incorporated environmental condition sensors.
@@
Although soft robotics is a promising research area with extensive applicability, it entails considerable challenges. One of the most intricate challenges is finding a proper design: finding the most suitable robot design requires substantial time and material resources because it implies numerous iterations of testing the design in real life [53]. Unfortunately, for soft robots, design and other aspects are more complex, since they use flexible materials whose mechanical properties are non-linear and complex to characterise [19]. For instance, biohybrid machines, a particular type of soft robots, are built employing biological components, such as cells or tissues, which are inherently difficult to control with conventional methods. This brings another set of challenges, since several parameters need to be considered during the design of these soft robots.
@@
One promising strategy to counteract the intrinsic challenges of soft robots is the use of NeuroEvolution (NE). NE consists of evolving the topology and weights of artificial neural networks (ANNs) through a genetic algorithm (GA). One of the most effective and well-known approaches is NeuroEvolution of Augmenting Topologies (NEAT) [57], which has demonstrated an adequate performance regarding morphology generation [7]. Moreover, NEAT was extended under the rationale that natural structures are composed of patterns and shape repetition. This extension is called Hypercube-based NeuroEvolution of Augmenting Topologies (HyperNEAT) [56]. This new approach evolves a particular type of networks, i.e., Compositional Pattern-Producing Networks (CPPNs), which use, among others, periodic functions, such as sine and cosine, to generate symmetry and patterns that help to evolve topologies [55].
@@
The primary objective of this research is to validate the suitability of NEAT and HyperNEAT as tools to design soft robot morphologies. Particularly an example is studied here of designing soft actuators of catheters for targeted drug delivery to areas of the human body that are difficult to reach. We chose the catheter scenario as an exemplary case study for automated morphology generation due to the simplicity and focus of its design. Catheters have limited degrees of freedom, so, the most critical factor in their morphology is bending. This streamlined focus on bending allows automated systems to concentrate on optimising flexibility, ensuring that generated models are both functional and adaptable to various medical scenarios.
@@
Furthermore, the presented research can be considered the continuation of the study presented in [4], where, in terms of designing soft robots for locomotion-oriented tasks, the suitability of NEAT and HyperNEAT was tested. The validation methodology implemented in this research compares the performance of NEAT and HyperNEAT against Age-Fitness Pareto Optimisation (AFPO), an evolutionary algorithm designed to avoid premature convergence [21]. Three metrics are used to assess the performance of the algorithms: (i) the general performance in terms of the displacement observed in one end of the actuators during a specific time-frame; (ii) the robustness of the designs of actuators produced; and (iii) the total volume of actuators.
@@
The rest of the paper is organised as follows. Section 2 describes background research focusing on soft-robot-related tasks utilising NEAT and HyperNEAT algorithms. Section 3 introduces in detail the mechanisms of the three algorithms utilised for experimentation in this research. Section 4 describes the experiments performed in the task of designing soft actuators for catheters. Finally, Section 5 concludes this work by presenting insights that emerged from the research and outlines future work.
@@ 2 Background
NEAT has been implemented in numerous studies focusing on soft-robot-oriented tasks. An example of these studies presents a three-link planar arm model driven by nine muscles, whose purpose is to simulate the human arm [63]. A neurocontroller trained by NEAT controls the musculor-skeleton arm. The training sets applied to NEAT are built considering the forward kinematics, geometry relationships, and muscle mechanic equations. NEAT is compared against a fixed-topology ANN. Results indicate that the neurocontroller trained by NEAT outperforms the traditional ANN in critical aspects such as moving the arm to a specific position and converging the distance between the endpoint and the target point with a minimum distance error.
@@
NEAT has not only been studied for controlling bioinspired mechanisms. It also has been utilised as a design engine for soft robots. For instance, in [11], NEAT, alongside CPPNs, is utilised to design soft robots capable of reaching or squeezing through a small aperture. The experiments consisted of simulations where each soft robot is placed within a cage whose dimensions are $15\times 15\times 11$, leaving a gap of 1 voxel (arbitrary volume in the physics simulator) in the $x$ and $y$ axes between the edge of the cage and the maximum size of the soft robot: $11\times 11\times 1$ 11. The cage is rigid, immobile and indestructible. Each side has an opening, approximately a circle of diameter 10. Although this research is primarily considered as a proof-of-concept, it is demonstrated that more compliant and deformable soft robots are more capable of reaching or squeezing through small apertures, such as tunnels, than soft robots that are less flexible.
@@
Following the same path of being utilized as morphology generator, in [6], NEAT (and CPPNs) is used to produce three-dimensional physical structures capable of conserving momentum to achieve maximum displacement due to gravity. Structures are composed of spherical cells which fuse to make rigid bodies. The growth process commences with a single cell called root and is located at the origin of the structure. A cloud of $n$ points is set around the cell with the $n$ points distributed on the root’s surface. Once the cloud is over the cell, every point is used to query a CPPN whose output represents the concentration of matter at that point. A matter threshold determines where to allocate a cell: the more the output value exceeds the matter threshold, the denser the cell allocated at that point will be. Results suggest that the proposed method can produce artefacts that suitably capture the non-obvious relationship between function and physical structure.
@@
On the other hand, HyperNEAT has been primarily utilised in tasks targeting the control of robot morphologies. An example of these studies presents a method where morphologies are the input, and the output is neural network controllers capable of working on these different morphologies [48]. Each neural controller is tested in three different quadruped morphologies whose lengths are 0.25, 0.30, and 0.37 meters. The neural controllers generated by HyperNEAT are compared against static controllers. Results indicate that HyperNEAT can identify the relationship between morphology and controller architectures during evolution, which leads to a suitable performance.
@@
In the same direction of controller design, in [12], a study that refers to legged robots, particularly four-legged, is presented. The rectangular torso of the robot morphology is 0.15 wide, 0.3 long, and 0.5 tall (in arbitrary ODE physics simulator units). Each leg has three cylinders (radius 0.02, length 0.075) and three hinge joints. HyperNEAT is compared against Fixed-Topology NEAT (FT-NEAT), also known as Perceptron NEAT. Results indicate that HyperNEAT outperforms FT-NEAT since its morphologies: (i) vastly outperform those generated by FT-NEAT in every generation and (ii) exhibit better coordination than those created by FT-NEAT. Another relevant insight of this research is that generative encodings can outperform direct encodings on regular problems.
@@
Another study, described in [66], focuses on creating gaits for quadruped robots. HyperNEAT is compared against six parameterised learning strategies: (a) uniform random hill climbing; (b) Gaussian random hill climbing; (c) policy gradient reinforcement learning; (d) Nelder-Mead simplex; (e) a random baseline; and (f) a new method that builds a model of the fitness landscape with linear regression. Nine servos actuate the robot used for experimentation, each commanded to a position in the $[0,1023]$ range, corresponding to a physical range of $[-120^{\circ},+120^{\circ}]$. Two on-board batteries power the servos. Results advocate the superiority of HyperNEAT over all parameterised learning strategies since the gaits it generates exhibit significantly better performance.
@@
Not only one neurocontroller can be evolved through HyperNEAT. In [14], it is stated that it is possible to evolve neurocontrollers encoded by a single genome representing a team of predator agents that work together to capture prey. In the experiments conducted, agents are not aware of their teammates, and due to prey avoidance from nearby predators, it is feasible for one predator to undermine another’s pursuit by knocking its prey off its path. Consequently, predators are forced to learn coherent roles that complement the roles of their allies. Results indicate that encoding a team of agents as a unified genome brings benefits such as: (i) critical skills do not need to be rediscovered for separate agents; (ii) due to CPPNs representing multiagent policies, they are assigned to separate agents as a function of their relative geometry and simultaneously exploiting the internal geometries of agents.
@@
More recently, another study tests the suitability of HyperNEAT (and NEAT) as a design engine for soft robot morphologies, which is described in [4]. The experiments focused on locomotion tasks, particularly reaching the maximum displacement possible in 10 seconds. Morphologies generated by AFPO are utilised as a baseline. Results indicate that neuroevolution-based approaches can outperform AFPO, and HyperNEAT can perform more suitably than NEAT, despite the absence of geometrical aspects of the problem domain in the design of the substrate. This confirms what is argued in [13]: HyperNEAT can obtain suitable results, even if there is no defined geometrical representation of the problem.
Thus, based on the previously analysed research findings, it is feasible to consider NEAT and HyperNEAT as the core strategy to design adequate morphologies for soft actuators for catheters.
3 Algorithms

### v2

@@ Abstract
Soft robots can exhibit better performance in specific tasks compared to conventional robots, particularly in healthcare related tasks. However, the field of soft robotics is still young, and designing them often involves mimicking natural organisms or relying heavily on human experts’ creativity. A formal automated design process is required. The use of neuroevolution-based algorithms to automatically design initial sketches of soft actuators that can enable the movement of future medical devices, such as drug-delivering catheters, is proposed. The actuator morphologies discovered by algorithms like Age-Fitness Pareto Optimisation, NeuroEvolution of Augmenting Topologies (NEAT), and Hypercube-based NEAT (HyperNEAT) were compared based on the maximum displacement reached and their robustness against various control methods. Analyzing the results granted the insight that neuroevolution-based algorithms produce better-performing and more robust actuators under diverse control methods. Specifically, the best-performing morphologies were discovered by the NEAT algorithm.
@@ 1 Introduction
Although soft robotics is a promising research area with extensive applicability, it entails considerable challenges. One of the most intricate challenges is discovering a proper design: finding the most suitable robot design requires substantial time and material resources, because, it implies numerous prototypes being tested in real life [64]. Unfortunately, for design of soft robots this is not the only complex aspect, since they use flexible materials whose mechanical properties are non-linear and complex to characterize [23]. For instance, biohybrid machines, a particular type of soft robots, are built employing biological components, such as cells or tissues, which are inherently difficult to control with conventional methods. This brings another set of challenges, since several parameters need to be considered during the design process and the reality gap problem is becoming more profound.
@@
One promising strategy to counteract these intrinsic challenges of soft robots is using NeuroEvolution (NE). This methodology was previously utilised to design reconfigurable biohybrid machines without strict definitions of specific targets [34]. NE consists of evolving the topology and weights of artificial neural networks (ANNs) through a genetic algorithm (GA). One of the most effective and well-known approaches is NeuroEvolution of Augmenting Topologies (NEAT) [70], which has demonstrated an adequate performance regarding morphology generation [8]. Moreover, NEAT was extended under the rationale that natural structures are composed of patterns and shape repetition. This extension is called Hypercube-based NeuroEvolution of Augmenting Topologies (HyperNEAT) [69]. This approach evolves a particular type of networks, i.e., Compositional Pattern-Producing Networks (CPPNs), which use, among others, periodic functions, such as sine and cosine, to generate symmetry and patterns that help to evolve topologies [68].
@@
The primary objective of this research is to design from scratch soft actuators that can be implemented in medical devices, such as catheters for targeted drug delivery in areas of the human body that are otherwise difficult to reach. In order to define a trustworthy design engine, NEAT and HyperNEAT are validated as tools to design soft actuator morphologies. We chose the catheter scenario as an exemplary case study for automated morphology generation due to its straightforward design. The inputs for the proposed methodology are the physical characteristics of two different types of material used as building blocks and the constraints that the actuator morphology needs to follow to resemble a catheter actuator. The fitness function of the possible solutions considered is the maximum bending angle that the actuator morphology can reach in silico. The output of the evolutionary methodologies is the best combination of the two predefined material building blocks that are composed together inside the predefined design space to produce the actuator morphology. This streamlined focus on bending, which allows other automated systems to concentrate on optimising flexibility, ensuring that generated models are both functional and adaptable to various medical scenarios.
NE algorithms were utilised here to evolve the design of soft actuators morphologies under an indirect representation of the three-dimensional space. Specifically, instead of using a bitstream of all possible dimensional segments to encode the characteristics of every point in the available design space (i.e. the type of material used for that point), the NE evolves the artificial networks that when queried with all the points in the design space, they return the material characteristic for each point of the morphology design.
@@
The validation methodology implemented in this research compares the performance of NEAT and HyperNEAT against Age-Fitness Pareto Optimisation (AFPO), an evolutionary algorithm designed to avoid premature convergence [25]. Three metrics are used to assess the performance of the algorithms: (i) the general performance in terms of the displacement observed in the free end of the actuator during a specific time-frame; (ii) the robustness of the designs of actuators produced; and (iii) the total volume of actuators. The fitness of the possible solutions is provided by a physics engine simulating the actuator. The aforementioned metrics and an updated set of constraints applied in the physics engine, present an altered search space for the algorithms which proved to be efficient in previous works of designing soft and biohybrid robots tested on purely the longer distance covered in unconstrained locomotion tasks [4, 34].
Whereas the methodology of NE in optimisation of morphologies of soft robots and biohybrid machines was previously studied [34, 71, 40], the contribution of this work is towards targeting a specific real-world application. Namely, previous works used evolution towards unconstrained locomotion with a generic fitness function (i.e., the longest point reached by a morphology and some variations of that but without any constraints), the present study describe an environment with specific constraints, in order to better reflect the application under study. In specific, here one end of the morphology is fixed, to simulate the insertion point of a drug-delivery catheter. Moreover, while previous works [34] used AFPO with mutation only operators, here we use NEAT and HyperNEAT with mutation and recombination operators to explore in more efficient way the design space of the problem.
@@
The rest of the paper is organized as follows. Section 2 describes background research focusing on soft robot related tasks and in specific utilising NEAT and HyperNEAT algorithms. Section 3 introduces in detail the automated design methodology and the mechanisms of the three algorithms utilised for experimentation in this research. Section 4 describes the experiments performed in the task of designing soft actuators for catheters. Finally, Section 5 concludes this work by presenting insights that emerged from the research and outlines future work.
@@ 2 Background
The concept of soft robots can be traced back to Isaac Asimov’s books published in the 1940s-1950s [6]. However, the first scientific and engineering papers related to soft grippers [24] and soft architecture machines [49] started in the late 1970s. The domain has started to flourish in late 1980s with further publications on soft fingers [3, 11, 12], electrostatic actuators [20], and fluid based grippers [31]. Unlike traditional and rigid robots, these robots are built utilising flexible and ductile materials [38, 61]. Soft robots have the potential to exhibit better performance in specific tasks that are related to search and rescue [46, 48, 76, 78], space exploration [45, 5, 50, 47], and healthcare [54, 39, 26] due to their mechanisms and morphology, which are inspired by the movement and behavior of living organisms.
Publications on soft robotic catheters, and relevant surgical devices, started to emerge from late 1980s [35] and by 2020 there was a range of catheters developed for cardio-vascular surgery [52, 60, 51], urinary system [36, 9, 79], endovascular treatment [22], in situ bioprinting [82], endoscopy [37] and drug delivery [27]. Plantoid robots proposed by Mazzolai and colleagues [42, 44, 19, 43, 41, 62] can be considered relevant to the research on robotic catheters. These robots are soft-bodied mechanical devices capable of physically growing at the tip (thanks to embedded 3D printing techniques), similar to a plant root, and can change the direction of their growth due to incorporated environmental condition sensors.
The use of soft robots in medical education is also noteworthy. For instance, a recent work [56] implements McKibben artificial muscle on a soft robot inspired by the large intestine and demonstrates similar behavior to its biological counterpart. This soft robot was designed and 3D printed in a prototype that was tested along with a simulation of finite element analysis to validate its performance. As a result, this soft robot may be employed by medical students to improve their understanding of the functionality of large intestine via optical interaction and touch. Moreover, the same robotic architecture was considered with an updated control mode to achieve higher speeds and loads of object interaction [55]. Drawing inspiration from the movement of inchworm, rather than the intestine peristalsis, enhanced its transport capabilities. The investigation of several of the design and control aspects provided significant insights into the enhanced efficiency and adaptability of its application on object transport tasks.
Several research studies have been conducted to find suitable strategies for soft robot design. For instance, in [10], a large-scale benchmark platform for optimising the design and control of soft robots is introduced. The software is called Evolution Gym, and robots can be composed of different types of voxels such as soft, rigid actuators. The benchmark environments contained in Evolution Gym provide a wide range of tasks, such as locomotion and object manipulation. Robots are designed in a two-dimensional layout.
Furthermore, another design tool for multi-material soft actuators, SoRoForge, is introduced in [66]. This platform comprises the major stages of this type of actuator’s design, evaluation, and fabrication process. The software allows: (i) an interactive exploration of implicit geometry functions represented by computational networks; (ii) seeing an online three-dimensional static geometry preview of designs; and (iii) the creation of three-dimensional printable design files.
Another design automation system for soft actuators is presented in [65]. The design engine is the well-known multi-objective evolutionary approach called Non-dominated Sorting Genetic Algorithm-II (NSGA-II) [18] that was implemented to find suitable morphologies of soft actuators. Results suggest that under this approach, it is necessary to define adequate fitness functions to find suitable morphologies of soft robots since the ones found in the series of experiments described in this work did not outperform those solutions that were a priori manually designed and were used as a baseline.
Evolutionary approaches are not the only methods employed to design soft robots. For instance, in [40], a gradient-based approach, called novo optimisation, is utilised to find suitable structures of robots. The algorithm is tested using terrestrial locomotion tasks and shown to be capable of: (i) assessing the fitness of the behavior of the soft robot; (ii) identifying deficiencies in its overall shape, topology, number and shape of limbs, mass distribution, musculature, and behavioral control; and (iii) simultaneously changing all these aspects which leads to an improvement in evaluation.
In the same path of non-evolutionary approaches, a study introduces a soft robot design methodology based on topology optimisation [81]. The authors implement their methodology to design a soft gripper capable of undergoing free travel and delivering a grasping force. Moreover, the design can be printed using three-dimensional additive manufacturing. However, the optimisation model assumes the homogeneity of materials.
@@
Regarding the two neuroevolution-based algorithms studied here, NEAT has been implemented in numerous studies focusing on soft robots. An example of these studies presents a three-link planar arm model driven by nine muscles, whose purpose is to simulate the human arm [77]. A neurocontroller trained by NEAT controls the muscular-skeleton arm. The training sets applied to NEAT are built considering the forward kinematics, geometry relationships, and muscle mechanic equations. NEAT is compared against a fixed-topology ANN. Results indicate that the neurocontroller trained by NEAT outperforms the traditional ANN in critical aspects such as moving the arm to a specific position and converging the distance between the endpoint and the target point with a minimum distance error.
@@
NEAT has not only been studied for controlling bioinspired mechanisms. It also has been utilised as a design engine for soft robots. For instance, in [13], NEAT, alongside CPPNs, is utilised to design soft robots capable of reaching or squeezing through a small aperture. The experiments consisted of simulations where each soft robot is placed within a cage whose dimensions are $15\times 15\times 11$ voxels (arbitrary volume in the physics simulator), leaving a gap of 1 voxel in the $x$ and $y$ axes between the edge of the cage and the maximum size of the soft robot: $11\times 11\times 1$ 11. The cage is rigid, immobile and indestructible. Each side has an opening, approximately a circle of diameter 10. Although this research is primarily considered as a proof-of-concept, it is demonstrated that more compliant and deformable soft robots are more capable of reaching or squeezing through small apertures, such as tunnels, than soft robots that are less flexible.
@@
Following the same path of being utilised as morphology generator, in [7], NEAT (and CPPNs) is used to produce three-dimensional physical structures capable of conserving momentum to achieve maximum displacement due to gravity. Structures are composed of spherical cells which fuse to make rigid bodies. The growth process commences with a single cell called root and is located at the origin of the structure. A cloud of $n$ points is set around the cell with the $n$ points distributed on the root’s surface. Once the cloud is over the cell, every point is used to query a CPPN whose output represents the concentration of matter at that point. A matter threshold determines where to allocate a cell: the more the output value exceeds the matter threshold, the denser the cell allocated at that point will be. Results suggest that the proposed method can produce artefacts that suitably capture the non-obvious relationship between function and physical structure.
@@
On the other hand, HyperNEAT has been primarily utilised in tasks targeting the control of robots. An example of these studies presents a method where morphologies are the input, and the output is neural network controllers capable of working on these different morphologies [59]. Each neural controller is tested in three different quadruped morphologies whose lengths are 0.25, 0.30, and 0.37 meters. The neural controllers generated by HyperNEAT are compared against static controllers. Results indicate that HyperNEAT can identify the relationship between morphology and controller architectures during evolution, which leads to a suitable performance.
@@
In the same direction of controller design, in [15] legged robots are studied, and particularly four-legged robots. The rectangular torso of the robot morphology is 0.15 wide, 0.3 long, and 0.5 tall (in arbitrary ODE physics simulator units). Each leg has three cylinders (radius 0.02, length 0.075) and three hinge joints. HyperNEAT is compared against Fixed-Topology NEAT (FT-NEAT), also known as Perceptron NEAT. Results indicate that HyperNEAT outperforms FT-NEAT since its morphologies: (i) vastly outperform those generated by FT-NEAT in every generation and (ii) exhibit better coordination than those created by FT-NEAT. Another relevant insight of this research is that generative encodings can outperform direct encodings on regular problems.
@@
Another study, described in [80], focuses on creating gaits for quadruped robots. HyperNEAT is compared against six parameterized learning strategies: (a) uniform random hill climbing; (b) Gaussian random hill climbing; (c) policy gradient reinforcement learning; (d) Nelder-Mead simplex; (e) a random baseline; and (f) a new method that builds a model of the fitness landscape with linear regression. Nine servos actuate the robot used for experimentation, each commanded to a position in the $[0,1023]$ range, corresponding to a physical range of $[-120^{\circ},+120^{\circ}]$. Two on-board batteries power the servos. Results advocate the superiority of HyperNEAT over all parameterized learning strategies since the gaits it generates exhibit significantly better performance.
@@
Not only standalone neurocontrollers can be evolved through HyperNEAT. In [17], it is stated that it is possible to evolve neurocontrollers encoded by a single genome representing a team of predator agents that work together to capture prey. In the experiments conducted, agents are not aware of their teammates, and due to prey avoidance from nearby predators, it is feasible for one predator to undermine another’s pursuit by knocking its prey off its path. Consequently, predators are forced to learn coherent roles that complement the roles of their allies. Results indicate that encoding a team of agents as a unified genome brings benefits such as: (i) critical skills do not need to be rediscovered for separate agents; (ii) due to CPPNs representing multiagent policies, they are assigned to separate agents as a function of their relative geometry and simultaneously exploiting the internal geometries of agents.
@@
More recently, another study tests the suitability of HyperNEAT (and NEAT) as a design engine for soft robot morphologies, which is described in [4]. The experiments focused on locomotion tasks, particularly reaching the maximum displacement possible in 10 seconds. Morphologies generated by AFPO are considered as a baseline. Results indicate that both neuroevolution-based approaches can outperform AFPO, and HyperNEAT can perform more suitably than NEAT, despite the absence of geometrical aspects of the problem domain in the design of the substrate. This confirms what is argued in [16]: HyperNEAT can obtain suitable results, even if there is no defined geometrical representation of the problem.
Thus, based on the aforementioned research findings, it is feasible to consider NEAT and HyperNEAT as the core strategy to design adequate morphologies for soft actuators for catheters. Furthermore, considering the three-dimensional nature of the morphologies, Voxelyze is a suitable software to simulate their physical behavior. Although Voxelyze has limited capacity in capturing complicated physical properties (i.e., does not take into account fluid dynamics), it was selected as an efficient alternative against more complicated simulators. The basic functionality (i.e., interactions between adjacent voxels, volumetric actuation and collision detection) is considered adequate under the trade off between a lower reality gap and a computational efficient evolutionary optimisation. Nonetheless, all candidate solutions in the evolutionary optimisation will undergo the same inaccuracies imposed by the simulator. Moreover, Voxelyze has demonstrated adequate accuracy in simulating biohybrid machines [34] that exhibit inherently high levels of behavioral uncertainty. This has been achieved with only one round of fine-tuning based on in vitro observations, aligning simulations more closely with physical prototypes.
In the medical domain, however, the accuracy required for final designs is paramount, as high fidelity in biohybrid actuator performance is essential. Therefore, before advancing any selected design to manufacturing prototyping, these preliminary designs should undergo a secondary stage of rigorous testing in simulations that account for more precise physics-based properties, including fluid dynamics and nuanced biological interactions. By coupling our approach with a more detailed simulator at next developmental stage, we can better approximate real-world performance, mitigating risks associated with physical prototyping and ensuring that any final designs meet the strict reliability and performance standards needed for medical applications.
Some previous studies focus in a two-dimensional layout design of soft robots [10, 71], the present work aims towards more realistic three-dimensional designs, thus, Voxelyze is utilised. Whereas some non-evolutionary approaches were previously proposed for soft robot design [40, 81], the utilisation of evolutionary methods is preferred here, due to their simplicity and enhanced capacity to avoid local optima in rugged optimisation problems. Nonetheless, the direct representation of individuals in evolutionary soft robot design techniques has proved not as efficient as generative representation (i.e. CPPNs) [14]. The implementation of AFPO has proved very efficient when employed in the designing process of biohybrid machines [34]. However, it was previously applied using only a mutation operator, without crossover, which limited the method’s exploratory capacity. The authors of the prior study recommended incorporating recombination-based methods, such as NEAT and HyperNEAT, for future improvements. Finally, the implementation of NE algorithms have been tested in specific applications, namely achieving the longest distance on an unconstrained plane or with some obstacles located on the plane and object manipulation [14, 34, 71, 40]. Here, we study the application of NE algorithms on a more reserved and specialized application (namely, the angular movement of a soft actuator enabling a medical catheter).
3 Methodology
The design of soft actuators in a given three-dimensional space ($[x,y,z]$) was produced automatically via an evolutionary process over an indirect representation and evaluated through a physics engine (i.e. Voxelyze) to fit the initially set target. The target (or the fitness function) here was set to achieve the higher possible bending angle, since, the application target is a drug-delivering catheter. To represent the insertion point of the catheter, one end of the morphology was fixed and not able to move, thus, emulated an anchoring point for the actuator. Moreover, instead of using a direct representation in the evolutionary optimisation for the design of the soft actuator, namely a one-to-one correspondence of all possible positions in the three-dimensional design space with their physical characteristics, an indirect representation was preferred. The indirect representation used in all of the algorithms described in the following, is constituted by neural networks, that have as input all the possible combination of the three-dimensional space ($[x,y,z]$) and as outputs the physical characteristics of the material comprising the actuator morphology in the corresponding area.
The fitness function consists of a physics engine called Voxelyze (also available in [2]), which simulates the physical response of morphologies under determined conditions. Morphologies are defined within Voxelyze by stacking voxels (the minimum building block in Voxelyze) of different materials together. Here, two types of voxels were considered for the simplicity of the designing landscape: one active and one passive voxel, which can represent the response of contractile muscles and immotile cells, respectively. The physics simulation parameters for both these building blocks were set to Young’s modulus of $5*10^{6}$ Pa, Poisson’s Ratio of $0.35$, and coefficients of static and dynamic friction of $1.0$ and $0.5$, respectively. Moreover, the active voxel is characterized by volumetric actuation of ±$\pm 50\%$ of the volume at rest at $4Hz$. These parameters were selected based on previously published and in vitro validated works [34], but do not limit the generalization of the algorithmic methodologies.
The original version of Voxelyze was modified (changes in the source code available in [1]) to trace the position of the catheter actuator in the $x$,$y$,$z$ axes during a time $t$. Furthermore, one end of actuators is fixed in such a way that they only move vertically (i.e., one degree of freedom) in the $yz$ plane. In addition, actuators have a passive enclosure to follow the findings of previous simulation works [74] and constraints of laboratory experiments, i.e. the bioreactor that a muscle actuator requires for the supply of nutrients in case of a biohybrid machine. The output of Voxelyze contains three values that describe the position of the actuator free tip (with respect to $x$,$y$,$z$ dimensions) for the initial position and the final one (after the predefined simulation time), and the number of voxels that compose the actuator.